# CtxMIM: Context-Enhanced Masked Image Modeling for Remote Sensing Image Understanding

### v3

@@ 4.1 Experimental Settings
Sup. We employ three general backbones (ResNet-50 (He et al., 2016), ViT (Dosovitskiy et al., 2020), and Swin Transformer (Liu et al., 2021)) initialized with ImageNet pretraining for supervised learning comparison. In experiments, “Random Init.” and “ImageNet Pre.” denote that different backbones are initialized with random initialization and ImageNet pretraining, respectively. Besides, SatLas (Bastani et al., 2023) is compared as a remote sensing supervised method since it is pretrained on a large-scale remote sensing dataset with 137 categories under seven label types.
@@
    SeCo (Manas et al., 2021), based on Momentum Contrast (MoCov2) (Chen et al., 2020b), pretrains a ResNet-50 backbone for 200 epochs with a batch size of 256. It collects a new dataset from Sentinel-2 (Drusch et al., 2012), designed to capture seasonal changes by sampling five images from distinct dates.
@@
    Geo-Aware (Ayush et al., 2021) pretrains a ResNet-50 backbone based on MoCov2 (Chen et al., 2020b) for 200 epochs with a batch size of 256. It creates GeoImageNet by extracting geo-coordinates for ImageNet (Russakovsky et al., 2015). Finally, Geo-Aware uses the Functional Map of the World (fMoW) (Christie et al., 2018) and GeoImageNet for pretraining.
@@
    CACo (Mall et al., 2023) adopts the same pretraining settings and backbone as SeCo (Manas et al., 2021). CACo reconsiders the temporal signal to contrast images with long-term and short-term differences, spanning four years for the long-term and one year for the short-term.
@@
    TOV (Tao et al., 2023) is the original vision model with a ResNet50 backbone in the remote sensing field. By analyzing the influences of different data sampling and the selection of learning paths, it releases a benchmark dataset for pretraining. Finally, TOV is pretrained for 800 epochs with a batch size of 1024.
@@
    CMID (Muhtar et al., 2023) is a unified SSL framework which learns representations by combining the CL and MIM. CMID pretrains ResNet-50 and Swin-B backbones for 200 epochs on the MillionAID dataset with a batch size of 512 and 256, respectively.
    Table 1: Top-1 accuracy on EuroSAT and NWPU-RESISC45 for land cover classification. CtxMIM obtains the best performance.
@@
    SatLas (Bastani et al., 2023) ResNet50 96.62 86.23
    SatLas (Bastani et al., 2023) Swin-B 83.44 79.88
    Self-Sup. SeCo (Manas et al., 2021) ResNet50 95.09 61.52
    Geo-Aware (Ayush et al., 2021) ResNet50 95.66 81.15
    CACo (Mall et al., 2023) ResNet50 96.59 68.60
    TOV (Tao et al., 2023) ResNet50 57.81 46.74
    CMID (Muhtar et al., 2023) ResNet50 91.07 76.07
    GeRSP (Huang et al., 2024) ResNet50 98.29 92.83
    ScaleMAE (Reed et al., 2023) ViT-L 86.12 78.04
    SatMAE++ (Noman et al., 2024) ViT-L 87.83 82.60
    CrossScaleMAE (Tang et al., 2024) ViT-L 72.81 84.76
    SpectralGPT (Hong et al., 2024) ViT-B 84.61 71.04
    GFM (Mendieta et al., 2023) Swin-B 88.87 84.42
    CMID (Muhtar et al., 2023) Swin-B 92.72 93.23
@@
  6. 6.
    GeRSP (Huang et al., 2024) is a MoCo-based framework using a teacher-student architecture. It pretrains a ResNet-50 backbone on the unlabeled Million-AID (MAID) (Long et al., 2021) and ImageNet (Russakovsky et al., 2015) for 200 epochs with a batch size of 128.
  7. 7.
    ScaleMAE (Reed et al., 2023) is based on the MAE architecture (He et al., 2022a), employing ViT-Large as the backbone. It uses the Functional Map of the World (FMoW) (Christie et al., 2018) RGB training set for 800 epochs with a batch size of 128.
  8. 8.
    SatMAE++ (Noman et al., 2024) is a multi-scale pre-training approach, employing a Mask Autoencoder (MAE) framework. SatMAE++ pretrains a ViT-Large (Dosovitskiy et al., 2020) model on the Functional Map of the World (FMoW) (Christie et al., 2018) RGB training set. It lasts for 800 epochs with a batch size of 64.
@@
    CrossScaleMAE (Tang et al., 2024), based on Masked Auto-Encoder (MAE), pretrains a ViT-Large backbone on the Functional Map of the World (fMoW) (Christie et al., 2018) RGB training set. It is pretrained for 400 epochs with a batch size of 512.
@@
    SpectralGPT (Hong et al., 2024) is a foundation model by considering unique characteristics of spectral data in a MAE framework. It pretrains a ViT-Base backbone on an extensive dataset from the Sentinel2 satellite with over one million images. Finally, SpectralGPT is pretrained for 200 epochs with a batch size of 16.
@@
    GFM (Mendieta et al., 2023) collects a compact yet diverse dataset GeoPile from multiple sources to promote feature diversity. GFM introduces a multi-objective continual pretraining paradigm based on MIM (Xie et al., 2022), which can leverage the representations of ImageNet to learn valuable features. Finally, it pretrains a Swin-B backbone for 100 epochs with a batch size of 128.
Table 2: Mean intersection over union (mIoU) and mean accuracy (mAcc) on SpaceNet (Rio) semantic segmentation task. Our method consistently improves supervised and self-supervised learning by large margins.
@@
SatLas (Bastani et al., 2023) ResNet50 74.56 82.38
SatLas (Bastani et al., 2023) Swin-B 67.86 76.22
Self-Sup. SeCo (Manas et al., 2021) ResNet50 73.89 81.63
Geo-Aware (Ayush et al., 2021) ResNet50 74.99 82.30
CACo (Mall et al., 2023) ResNet50 74.98 82.77
TOV (Tao et al., 2023) ResNet50 74.12 81.89
CMID (Muhtar et al., 2023) ResNet50 71.53 79.28
GeRSP (Huang et al., 2024) ResNet50 76.29 84.47
ScaleMAE (Reed et al., 2023) ViT-L 77.21 85.12
SatMAE++ (Noman et al., 2024) ViT-L 77.25 85.40
CrossScaleMAE (Tang et al., 2024) ViT-L 75.55 83.87
SpectralGPT (Hong et al., 2024) ViT-B 69.17 77.78
GFM (Mendieta et al., 2023) Swin-B 72.81 81.71
CMID (Muhtar et al., 2023) Swin-B 65.98 73.72
@@
4.2 Land Cover Classification
Implementation Details. We take two land cover classification datasets: (1) EuroSAT (Helber et al., 2019) is captured by the Sentinel-2 satellite over European cities. This dataset has 27,000 labeled images with a size of 64$\times$64. It contains 10 classes with 2,000-3,000 images per class. Following SeCo (Manas et al., 2021), we adopt the official train/val splits on EuroSAT. (2) NWPU-RESISC45 (Cheng et al., 2017) is a publicly available benchmark for remote sensing classification, which is published by Northwestern Polytechnical University (NWPU). The dataset contains 31,500 images with spatial resolutions ranging from 30 to 0.2m and a size of 256$\times$256. It covers 45 categories with 700 images in each category. Following GeRSP (Huang et al., 2024), we use the same train/val splits for this experiment.
Quantitative Results. Table 1 shows the top-1 accuracy of different methods on two datasets using fully supervised and self-supervised learning. For the EuroSAT dataset, we can see that CtxMIM outperforms the ImageNet supervised Swin-B backbone by +5.49%, illustrating that SSL on a large-scale remote sensing dataset improves performance for remote sensing downstream tasks. Additionally, CtxMIM achieves +6.15% higher top-1 accuracy than reconstructive learning methods, demonstrating that our approach learns robust and good transferable remote sensing representations for the land cover classification task. For the NWPU-RESISC45 dataset, we observe that CtxMIM improves the model of the Swin-B backbone initialized with the ImageNet supervised pretraining by +4.4% higher, proving that the domain gap exists between natural and remote sensing images. We find that CtxMIM outperforms contrastive and reconstructive pretraining methods by +2.37% and +1.97% higher, respectively.
Table 3: Average precision (AP) and average recall (AR) under different intersection over union thresholds on DOTA object detection task. $F1$/$F1_{75}$ calculated by $AP$/$AP_{75}$ and $AR$/$AR_{75}$ measures the performance more comprehensively/accurately.
@@ 4.2 Land Cover Classification
SatLas (Bastani et al., 2023) ResNet50 30.0 52.8 31.5 15.8 30.7 38.7 41.1 69.4 43.7 22.9 44.0 54.0 34.7 36.6
SatLas (Bastani et al., 2023) Swin-B 27.8 50.7 28.7 18.0 28.6 34.0 42.5 75.1 43.8 27.7 46.9 52.7 33.6 34.7
SeCo (Manas et al., 2021) ResNet50 32.0 55.4 33.9 18.6 32.7 38.5 44.7 74.7 47.6 28.0 48.1 55.1 37.3 39.6
Geo-Aware (Ayush et al., 2021) ResNet50 33.2 58.0 34.3 17.8 34.4 41.9 44.5 73.9 47.4 26.7 47.9 56.1 38.0 39.8
CACo (Mall et al., 2023) ResNet50 33.7 58.5 35.4 19.9 34.6 41.4 45.7 75.5 48.5 28.8 48.8 56.8 38.8 40.9
TOV (Tao et al., 2023) ResNet50 35.2 61.7 36.6 21.6 36.2 44.2 46.4 76.8 49.3 30.1 49.3 58.5 40.0 42.0
CMID (Muhtar et al., 2023) ResNet50 36.3 63.0 38.3 21.1 38.0 44.8 48.1 79.9 51.8 28.8 52.9 60.0 41.4 44.0
GeRSP (Huang et al., 2024) ResNet50 37.6 63.3 39.4 22.0 38.6 47.3 47.9 76.0 51.9 29.8 51.4 60.2 42.1 44.8
GFM (Mendieta et al., 2023) Swin-B 35.4 62.7 36.8 22.5 38.0 44.1 48.8 81.4 52.1 32.0 53.5 60.1 41.0 43.1
CMID (Muhtar et al., 2023) Swin-B 27.3 51.0 27.5 17.2 28.7 32.6 42.2 75.9 43.1 26.4 47.0 52.0 33.2 33.6
@@
4.3 Semantic Segmentation
Implementation Details. We use the SpaceNet (Rio) dataset (Van Etten et al., 2018) for the semantic segmentation task. This dataset contains 6,940 satellite images with binary building masks. We fine-tune a PSANet (Zhao et al., 2018) with the pretrained Swin-B backbone by CtxMIM. We use an AdamW optimizer with a batch size of 8 and a learning rate of $1\times 10^{-4}$. We report mean intersection over union (mIoU) and mean accuracy (mAcc) to compare CtxMIM with comparison methods.
Quantitative Results. Table 2 presents mIoU and mAcc values of different methods on SpaceNet (Rio). The results show that our method achieves more considerable performance gains by +6.66% and +5.99% for supervised learning from the ImageNet pretrained weight with a Swin-B backbone. Compared with self-supervised learning, our method improves the mIoU performance by +2.93% and +1.97% over SOTA contrastive and reconstructive learning methods. Consistent experimental results demonstrate the effectiveness of our method for the semantic segmentation task.
4.4 Object Detection
Implementation Details. We use the DOTA dataset (Xia et al., 2018) to verify the object detection performance. DOTA dataset consists of 2,806 high-resolution satellite images with sizes ranging from 800$\times$800 to 4,000$\times$4,000. This dataset has 15 different categories with bounding box annotations and is split into 800$\times$800 following GeRSP (Huang et al., 2024). Then, we fine-tune a RetinaNet (Lin et al., 2017) with an AdamW optimizer, a batch size of 4, a learning rate of $1\times 10^{-5}$, and a weight decay of 0.05. We use the standard MS-COCO metrics (Lin et al., 2014), including average precision (AP) and average recall (AR) under different intersection over union thresholds, to compare CtxMIM with supervised and self-supervised methods on this task. Moreover, we calculate $F1$ and $F1_{75}$ to evaluate CtxMIM more comprehensively and accurately.
Quantitative Results. Table 3 reports results from comprehensive experiments on the DOTA dataset compared with supervised and self-supervised learning methods. CtxMIM consistently obtains the best performance across supervised and self-supervised learning methods. Specifically, CtxMIM gains +1.4% and +3% improvements on $F1$ and $F1_{75}$ over ImageNet supervised learning with a Swin-B backbone. Furthermore, CtxMIM improves the indicator $F1$ by +1% and 2.1% compared with contrastive and reconstructive learning methods. The extraordinary performance demonstrates that CtxMIM learns impressive representations for the object detection task.
Table 4: $AP^{m}$, $AP_{50}^{m}$, and $AP_{75}^{m}$ on SpaceNet (Las Vegas) instance segmentation task. CtxMIM outperforms comparison methods (+2.2% and +1.1% on $AP^{m}_{75}$ for Sup. and Self-Sup., respectively).
@@ 4.4 Object Detection
SatLas (Bastani et al., 2023) ResNet50 39.3 73.4 39.3
SatLas (Bastani et al., 2023) Swin-B 29.8 59.2 27.3
Self-Sup. SeCo (Manas et al., 2021) ResNet50 40.8 73.7 42.4
Geo-Aware (Ayush et al., 2021) ResNet50 48.9 80.9 56.3
CACo (Mall et al., 2023) ResNet50 44.9 77.8 49.4
TOV (Tao et al., 2023) ResNet50 41.5 74.2 43.8
CMID (Muhtar et al., 2023) ResNet50 42.9 74.1 47.8
GeRSP (Huang et al., 2024) ResNet50 54.3 85.1 65.4
ScaleMAE (Reed et al., 2023) ViT-L 60.0 89.4 71.8
SatMAE++ (Noman et al., 2024) ViT-L 59.4 89.0 71.7
CrossScaleMAE (Tang et al., 2024) ViT-L 58.3 87.5 70.0
SpectralGPT (Hong et al., 2024) ViT-B 57.9 88.0 69.2
GFM (Mendieta et al., 2023) Swin-B 47.2 80.0 53.5
CMID (Muhtar et al., 2023) Swin-B 32.7 63.6 31.3
@@
Table 5: Ablation study. LcC, SS, OD, and IS denote land cover classification, semantic segmentation, object detection, and instance segmentation, respectively. ‘- C-E’ removes the context-enhanced generative branch from CtxMIM.
@@
4.5 Instance Segmentation
Implementation Details. SpaceNet (LasVegas) dataset (Van Etten et al., 2018) is used for instance segmentation task, consisting of 3,851 images across Las Vegas collected from WorldView-3 satellites. The size of images is 650$\times$650 with a spatial resolution of 30 cm/pixel. The dataset contains 151,367 building polygon footprints in GeoJSON format. In our experiment, this dataset is randomly divided into train/test/validation subsets with a ratio of 8:1:1. We fine-tune MaskRCNN (He et al., 2017) with the ed Swin-B backbone from our CtxMIM on SpaceNet (LasVegas), using an AdamW optimizer, a learning rate of 0.0001, a batch size of 4. Finally, we report $AP^{m}$, $AP_{50}^{m}$, and $AP_{75}^{m}$ under different intersections over union thresholds in the segmentation level.
Quantitative Results. As shown in Table 4, our approach learns transferable representations, outperforming all the comparison methods. Compared with supervised learning, CtxMIM improves the ImageNet supervised pretraining by +1.7% $AP^{m}$ and +2.2% $AP_{75}^{m}$, respectively. Moreover, CtxMIM achieves +0.5% $AP^{m}$ and +1.1% $AP_{75}^{m}$ higher compared with the SOTA self-supervised learning method ScaleMAE (Reed et al., 2023). The impressive results illustrate that CtxMIM learns good representations with high-level semantic and fine-grained details.
4.6 Ablation Study
We conduct ablation studies to verify the effectiveness of the designed context-enhanced generative branch on four typical remote sensing tasks, respectively. We adopt general datasets and critical indicators to measure performance, including EuroSAT (Helber et al., 2019), SpaceNet (Rio) (Van Etten et al., 2018), DOTA (Xia et al., 2018), and SpaceNet (LasVegas) (Van Etten et al., 2018). As reported in Table 5, CtxMIM outperforms CtxMIM without the context-enhanced generative branch by +1.97% on top-1 accuracy, 2.31% on mIoU, +2.7% on $F1_{75}$, and +1.4% on $AP^{m}$. The consistent improvements have confirmed that the context-enhanced generative branch can significantly improve the performance, illustrating its superiority.

### v4

@@ IV-A Experimental Settings
Sup. We employ three general backbones (ResNet-50 [57], ViT [58], and Swin Transformer [55]) initialized with ImageNet pretraining for supervised learning comparison. In experiments, “Random Init.” and “ImageNet Pre.” denote that different backbones are initialized with random initialization and ImageNet pretraining, respectively. Besides, SatLas [59] is compared as a remote sensing supervised method since it is pretrained on a large-scale remote sensing dataset with 137 categories under seven label types.
@@
    SeCo [3], based on Momentum Contrast (MoCov2) [6], pretrains a ResNet-50 backbone for 200 epochs with a batch size of 256. It collects a new dataset from Sentinel-2 [60], designed to capture seasonal changes by sampling five images from distinct dates.
@@
    Geo-Aware [4] pretrains a ResNet-50 backbone based on MoCov2 [6] for 200 epochs with a batch size of 256. It creates GeoImageNet by extracting geo-coordinates for ImageNet [1]. Finally, Geo-Aware uses the Functional Map of the World (fMoW) [61] and GeoImageNet for pretraining.
@@
    CACo [5] adopts the same pretraining settings and backbone as SeCo [3]. CACo reconsiders the temporal signal to contrast images with long-term and short-term differences, spanning four years for the long-term and one year for the short-term.
@@
    TOV [62] is the original vision model with a ResNet50 backbone in the remote sensing field. By analyzing the influences of different data sampling and the selection of learning paths, it releases a benchmark dataset for pretraining. Finally, TOV is pretrained for 800 epochs with a batch size of 1024.
@@
    CMID [63] is a unified SSL framework which learns representations by combining the CL and MIM. CMID pretrains ResNet-50 and Swin-B backbones for 200 epochs on the MillionAID dataset with a batch size of 512 and 256, respectively.
  6. 6.
    GeRSP [38] is a MoCo-based framework using a teacher-student architecture. It pretrains a ResNet-50 backbone on the unlabeled Million-AID (MAID) [64] and ImageNet [1] for 200 epochs with a batch size of 128.
  7. 7.
    ScaleMAE [44] is based on the MAE architecture [9], employing ViT-Large as the backbone. It uses the Functional Map of the World (FMoW) [61] RGB training set for 800 epochs with a batch size of 128.
  8. 8.
    SatMAE++ [43] is a multi-scale pre-training approach, employing a Mask Autoencoder (MAE) framework. SatMAE++ pretrains a ViT-Large [58] model on the Functional Map of the World (FMoW) [61] RGB training set. It lasts for 800 epochs with a batch size of 64.
    TABLE I: Top-1 accuracy on EuroSAT and NWPU-RESISC45 for land cover classification. CtxMIM obtains the best performance.
@@
    SatLas [59] ResNet50 96.62 86.23
    SatLas [59] Swin-B 83.44 79.88
    Self-Sup. SeCo [3] ResNet50 95.09 61.52
    Geo-Aware [4] ResNet50 95.66 81.15
    CACo [5] ResNet50 96.59 68.60
    TOV [62] ResNet50 57.81 46.74
    CMID [63] ResNet50 91.07 76.07
    GeRSP [38] ResNet50 98.29 92.83
    ScaleMAE [44] ViT-L 86.12 78.04
    SatMAE++ [43] ViT-L 87.83 82.60
    CrossScaleMAE [45] ViT-L 72.81 84.76
    SpectralGPT [42] ViT-B 84.61 71.04
    GFM [40] Swin-B 88.87 84.42
    CMID [63] Swin-B 92.72 93.23
@@
    CrossScaleMAE [45], based on Masked Auto-Encoder (MAE), pretrains a ViT-Large backbone on the Functional Map of the World (fMoW) [61] RGB training set. It is pretrained for 400 epochs with a batch size of 512.
@@
    SpectralGPT [42] is a foundation model by considering unique characteristics of spectral data in a MAE framework. It pretrains a ViT-Base backbone on an extensive dataset from the Sentinel2 satellite with over one million images. Finally, SpectralGPT is pretrained for 200 epochs with a batch size of 16.
@@
    GFM [40] collects a compact yet diverse dataset GeoPile from multiple sources to promote feature diversity. GFM introduces a multi-objective continual pretraining paradigm based on MIM [2], which can leverage the representations of ImageNet to learn valuable features. Finally, it pretrains a Swin-B backbone for 100 epochs with a batch size of 128.
IV-B Land Cover Classification
Implementation Details. We take two land cover classification datasets: (1) EuroSAT [10] is captured by the Sentinel-2 satellite over European cities. This dataset has 27,000 labeled images with a size of 64$\times$64. It contains 10 classes with 2,000-3,000 images per class. Following SeCo [3], we adopt the official train/val splits on EuroSAT. (2) NWPU-RESISC45 [11] is a publicly available benchmark for remote sensing classification, which is published by Northwestern Polytechnical University (NWPU). The dataset contains 31,500 images with spatial resolutions ranging from 30 to 0.2m and a size of 256$\times$256. It covers 45 categories with 700 images in each category. Following GeRSP [38], we use the same train/val splits for this experiment.
Quantitative Results. Table I shows the top-1 accuracy of different methods on two datasets using fully supervised and self-supervised learning. For the EuroSAT dataset, we can see that CtxMIM outperforms the ImageNet supervised Swin-B backbone by +5.49%, illustrating that SSL on a large-scale remote sensing dataset improves performance for remote sensing downstream tasks. Additionally, CtxMIM achieves +6.15% higher top-1 accuracy than reconstructive learning methods, demonstrating that our approach learns robust and good transferable remote sensing representations for the land cover classification task. For the NWPU-RESISC45 dataset, we observe that CtxMIM improves the model of the Swin-B backbone initialized with the ImageNet supervised pretraining by +4.4% higher, proving that the domain gap exists between natural and remote sensing images. We find that CtxMIM outperforms contrastive and reconstructive pretraining methods by +2.37% and +1.97% higher, respectively.
TABLE II: Mean intersection over union (mIoU) and mean accuracy (mAcc) on SpaceNet (Rio) semantic segmentation task. Our method consistently improves supervised and self-supervised learning by large margins.
@@ IV-B Land Cover Classification
SatLas [59] ResNet50 74.56 82.38
SatLas [59] Swin-B 67.86 76.22
Self-Sup. SeCo [3] ResNet50 73.89 81.63
Geo-Aware [4] ResNet50 74.99 82.30
CACo [5] ResNet50 74.98 82.77
TOV [62] ResNet50 74.12 81.89
CMID [63] ResNet50 71.53 79.28
GeRSP [38] ResNet50 76.29 84.47
ScaleMAE [44] ViT-L 77.21 85.12
SatMAE++ [43] ViT-L 77.25 85.40
CrossScaleMAE [45] ViT-L 75.55 83.87
SpectralGPT [42] ViT-B 69.17 77.78
GFM [40] Swin-B 72.81 81.71
CMID [63] Swin-B 65.98 73.72
@@
TABLE III: Average precision (AP) and average recall (AR) under different intersection over union thresholds on DOTA object detection task. $F1$/$F1_{75}$ calculated by $AP$/$AP_{75}$ and $AR$/$AR_{75}$ measures the performance more comprehensively/accurately.
@@
SatLas [59] ResNet50 30.0 52.8 31.5 15.8 30.7 38.7 41.1 69.4 43.7 22.9 44.0 54.0 34.7 36.6
SatLas [59] Swin-B 27.8 50.7 28.7 18.0 28.6 34.0 42.5 75.1 43.8 27.7 46.9 52.7 33.6 34.7
SeCo [3] ResNet50 32.0 55.4 33.9 18.6 32.7 38.5 44.7 74.7 47.6 28.0 48.1 55.1 37.3 39.6
Geo-Aware [4] ResNet50 33.2 58.0 34.3 17.8 34.4 41.9 44.5 73.9 47.4 26.7 47.9 56.1 38.0 39.8
CACo [5] ResNet50 33.7 58.5 35.4 19.9 34.6 41.4 45.7 75.5 48.5 28.8 48.8 56.8 38.8 40.9
TOV [62] ResNet50 35.2 61.7 36.6 21.6 36.2 44.2 46.4 76.8 49.3 30.1 49.3 58.5 40.0 42.0
CMID [63] ResNet50 36.3 63.0 38.3 21.1 38.0 44.8 48.1 79.9 51.8 28.8 52.9 60.0 41.4 44.0
GeRSP [38] ResNet50 37.6 63.3 39.4 22.0 38.6 47.3 47.9 76.0 51.9 29.8 51.4 60.2 42.1 44.8
GFM [40] Swin-B 35.4 62.7 36.8 22.5 38.0 44.1 48.8 81.4 52.1 32.0 53.5 60.1 41.0 43.1
CMID [63] Swin-B 27.3 51.0 27.5 17.2 28.7 32.6 42.2 75.9 43.1 26.4 47.0 52.0 33.2 33.6
@@
IV-C Semantic Segmentation
Implementation Details. We use the SpaceNet (Rio) dataset [13] for the semantic segmentation task. This dataset contains 6,940 satellite images with binary building masks. We fine-tune a PSANet [65] with the pretrained Swin-B backbone by CtxMIM. We use an AdamW optimizer with a batch size of 8 and a learning rate of $1\times 10^{-4}$. We report mean intersection over union (mIoU) and mean accuracy (mAcc) to compare CtxMIM with comparison methods.
Quantitative Results. Table II presents mIoU and mAcc values of different methods on SpaceNet (Rio). The results show that our method achieves more considerable performance gains by +6.66% and +5.99% for supervised learning from the ImageNet pretrained weight with a Swin-B backbone. Compared with self-supervised learning, our method improves the mIoU performance by +2.93% and +1.97% over SOTA contrastive and reconstructive learning methods. Consistent experimental results demonstrate the effectiveness of our method for the semantic segmentation task.
IV-D Object Detection
Implementation Details. We use the DOTA dataset [12] to verify the object detection performance. DOTA dataset consists of 2,806 high-resolution satellite images with sizes ranging from 800$\times$800 to 4,000$\times$4,000. This dataset has 15 different categories with bounding box annotations and is split into 800$\times$800 following GeRSP [38]. Then, we fine-tune a RetinaNet [66] with an AdamW optimizer, a batch size of 4, a learning rate of $1\times 10^{-5}$, and a weight decay of 0.05. We use the standard MS-COCO metrics [67], including average precision (AP) and average recall (AR) under different intersection over union thresholds, to compare CtxMIM with supervised and self-supervised methods on this task. Moreover, we calculate $F1$ and $F1_{75}$ to evaluate CtxMIM more comprehensively and accurately.
Quantitative Results. Table III reports results from comprehensive experiments on the DOTA dataset compared with supervised and self-supervised learning methods. CtxMIM consistently obtains the best performance across supervised and self-supervised learning methods. Specifically, CtxMIM gains +1.4% and +3% improvements on $F1$ and $F1_{75}$ over ImageNet supervised learning with a Swin-B backbone. Furthermore, CtxMIM improves the indicator $F1$ by +1% and 2.1% compared with contrastive and reconstructive learning methods. The extraordinary performance demonstrates that CtxMIM learns impressive representations for the object detection task.
TABLE IV: $AP^{m}$, $AP_{50}^{m}$, and $AP_{75}^{m}$ on SpaceNet (Las Vegas) instance segmentation task. CtxMIM outperforms comparison methods (+2.2% and +1.1% on $AP^{m}_{75}$ for Sup. and Self-Sup., respectively).
@@ IV-D Object Detection
SatLas [59] ResNet50 39.3 73.4 39.3
SatLas [59] Swin-B 29.8 59.2 27.3
Self-Sup. SeCo [3] ResNet50 40.8 73.7 42.4
Geo-Aware [4] ResNet50 48.9 80.9 56.3
CACo [5] ResNet50 44.9 77.8 49.4
TOV [62] ResNet50 41.5 74.2 43.8
CMID [63] ResNet50 42.9 74.1 47.8
GeRSP [38] ResNet50 54.3 85.1 65.4
ScaleMAE [44] ViT-L 60.0 89.4 71.8
SatMAE++ [43] ViT-L 59.4 89.0 71.7
CrossScaleMAE [45] ViT-L 58.3 87.5 70.0
SpectralGPT [42] ViT-B 57.9 88.0 69.2
GFM [40] Swin-B 47.2 80.0 53.5
CMID [63] Swin-B 32.7 63.6 31.3
@@
TABLE V: Ablation study. LcC, SS, OD, and IS denote land cover classification, semantic segmentation, object detection, and instance segmentation, respectively. ‘- C-E’ removes the context-enhanced generative branch from CtxMIM.
@@
IV-E Instance Segmentation
Implementation Details. SpaceNet (LasVegas) dataset [13] is used for instance segmentation task, consisting of 3,851 images across Las Vegas collected from WorldView-3 satellites. The size of images is 650$\times$650 with a spatial resolution of 30 cm/pixel. The dataset contains 151,367 building polygon footprints in GeoJSON format. In our experiment, this dataset is randomly divided into train/test/validation subsets with a ratio of 8:1:1. We fine-tune MaskRCNN [68] with the ed Swin-B backbone from our CtxMIM on SpaceNet (LasVegas), using an AdamW optimizer, a learning rate of 0.0001, a batch size of 4. Finally, we report $AP^{m}$, $AP_{50}^{m}$, and $AP_{75}^{m}$ under different intersections over union thresholds in the segmentation level.
Quantitative Results. As shown in Table IV, our approach learns transferable representations, outperforming all the comparison methods. Compared with supervised learning, CtxMIM improves the ImageNet supervised pretraining by +1.7% $AP^{m}$ and +2.2% $AP_{75}^{m}$, respectively. Moreover, CtxMIM achieves +0.5% $AP^{m}$ and +1.1% $AP_{75}^{m}$ higher compared with the SOTA self-supervised learning method ScaleMAE [44]. The impressive results illustrate that CtxMIM learns good representations with high-level semantic and fine-grained details.
Refer to captionRefer to caption
Figure 6: Some CtxMIM reconstructive samples on xView [69] and SpaceNet [13]. Each row shows the ground truth, four 2-tuples (the masked image and the CtxMIM reconstruction) of different masking ratios 70%, 75%, 80%, and 85% from left to right.
IV-F Ablation Study
We conduct ablation studies to verify the effectiveness of the designed context-enhanced generative branch on four typical remote sensing tasks, respectively. We adopt general datasets and critical indicators to measure performance, including EuroSAT [10], SpaceNet (Rio) [13], DOTA [12], and SpaceNet (LasVegas) [13]. As reported in Table V, CtxMIM outperforms CtxMIM without the context-enhanced generative branch by +1.97% on top-1 accuracy, 2.31% on mIoU, +2.7% on $F1_{75}$, and +1.4% on $AP^{m}$. The consistent improvements have confirmed that the context-enhanced generative branch can significantly improve the performance, illustrating its superiority.